# Fine-Grained Image-Text Alignment in Medical Imaging Enables Cyclic Image-Report Generation

### v4

@@ 5 Conclusion
We propose an AdaMatch for fine-grained image-text alignment, which presents the first work to adaptively associate image patches with words and improve the explainability of cyclic CXR-report generation. It includes the AdaPatch to acquire adaptive patches for abnormal regions, and performs patch-word alignment between adaptive patches with textual tokens. We implement cyclic CXR-report generation by using AdaMatch to provide explanations for the generation process. In addition, the fine-grained cyclic generation process supplies a natural explainability for the alignment between medical images and reports. Extensive experiments on two publicly available CXR datasets prove the effectiveness of our method and its superiority over previous methods.
Supplementary Material
In this supplementary material, we provide additional details to augment the content of the main paper. In Section 6, we conduct the ablation studies on the proposed AdaMatch-Cyclic. In Section 7, we demonstrate the visual results for CXR-to-report and report-to-CXR generation, and also visualize the adaptive patches and texts. In Section 8, we show implementation details of the proposed method.
6 Ablation Studies on AdaMatch-Cyclic
@@
7 Visual Results
Table 7: The effectiveness of keywords and keypatches.
Keywords Keypatches BLEU-4 METEOR ROUGE-L FID
@@ 7 Visual Results
Table 8: The analysis on the different number of keywords ($N_{w}$).
$N_{w}$ BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L
@@
Figure 11: Visualization of texts and the corresponding adaptive patches. The boxes and arrows in different colors show adaptive patches and their shift of center from fixed patches.
@@
CXR-to-Report Generation. As shown in Fig. 9, we compare CXR-to-report generation performance with existing methods on the MIMIC-CXR and OpenI datasets. In the first row, the proposed AdaMatch-Cyclic can capture the ‘Feeding tube tip’ and ‘mild bibasilar atelectasis’, while M2KT [47] cannot observe such device and lung abnormality. The second case of the OpenI dataset shows that ‘cardiac silhouette is mildly enlarged’ and ‘degenerative changes of the thoracic spine’ can be discovered by our method. These imply that our AdaMatch-Cyclic can generate a more comprehensive and complete medical report with the guidance of generated keywords in comparison to current methods.
@@
8 Implementation Details
In the AdaMatch-Cyclic model, we first train the AdaMatch model and then use the frozen AdaMatch to train LLM. The AdaMatch model consists of an image encoder and a text encoder. We utilize the PVT-medium [43] as the image encoder that includes AdaPatch module in stage 2, 3, and 4. The image encoder is pre-trained on the MIMIC-CXR dataset with the disease classification task. We adopt the pre-trained BioClinicalBERT [1] as text encoder. The image and text encoders are followed by two convolutional layers with batch normalization and the ReLU activation function, respectively, to reduce the feature dimension to 256. The patch size $s$ is set to 4 for stage 1, and 2 for stage 2, 3, and 4. The number of sampled feature points $m$ is 3. We optimize the AdaMatch using LAMB optimizer [51] ($\beta_{1}=0.9,\beta_{2}=0.999,\epsilon={10}^{-4}$). The cosine learning rate scheduler [29] with the base learning rate ($lr\_b$) of $6\times 10^{-3}$ is adopted to linearly warm up to the peak learning rate (i.e. $lr\_p=lr\_b\times\sqrt{\frac{bs\_all}{512}}$) during the first quarter of the total training epochs, where $bs\_all$ denotes the effective total batch size. The total training epoch is 15 and the per GPU batch size is 112. In AdaMatch-Cyclic, we use the pre-trained VQ-GAN [13] model to encode CXR images into image tokens and decode image tokens into CXR images. We adopt the dolly-v2-3b [9] model as the pre-trained LLM. The LLM has 5,1845 token types with the first 5,0821 token types for text tokens and the rest 1,024 token types for image tokens. We add 5,0821 to each image token value encoded by VQ-GAN. To decode the image tokens into images, we subtract 5,0821 from the image tokens generated by LLM and feed image tokens to VQ-GAN decoder to obtain the generated CXR images. We train the LLM with the AdamW [30] optimizer. The learning rate is initialized as $5\times{10}^{-6}$ and the total training epoch is 5. The per GPU batch size is set to 24. The hyper-parameters $\kappa_{0},\kappa_{1},\kappa_{2},\kappa_{3},\kappa_{4}$ are set as 200, 1, 1000, 20, and 5, respectively. In CXR-to-report generation, we extract keywords for each adaptive patch and use the keywords with the top 10 patch-word similarities as hints of LLM. In addition, we use 5 keypatches to guide the LLM in report-to-CXR generation. All the experiments are conducted on 8 Nvidia A100 40GB GPUs. The source code will be released.

### v5

@@ 5 Conclusion
We propose AdaMatch for fine-grained image-text alignment, which presents the first work to adaptively associate image patches with words and improve the explainability of cyclic CXR-report generation. It includes AdaPatch to acquire adaptive patches for abnormal regions and performs patch-word alignment between adaptive patches with textual tokens. We implement cyclic CXR-report generation by using AdaMatch to provide explanations for the generation process. In addition, the fine-grained cyclic generation process provides a natural explainability for the alignment between CXR images and reports. Extensive experiments on two CXR datasets show the effectiveness of our method and its superiority over previous methods.
Limitation
While our research has made significant strides in utilizing chest X-ray datasets, it is important to acknowledge certain limitations. Our experiments predominantly focus on chest X-ray datasets due to their availability of large-scale images paired with high-quality medical reports. However, these datasets primarily consist of patients from the ICU, potentially skewing our model towards severe disease domains. In our future endeavors, we intend to expand the scope of our methodology by applying the proposed AdaMatch-Cyclic to multi-domain scenarios, thereby mitigating this limitation and enhancing the versatility of our approach.
Acknowledge
This work was supported by the Hong Kong Research Grants Council (RGC) General Research Fund under Grant 14220622, Innovation and Technology Commission Innovation and Technology Fund ITS/229/22, and the National Natural Science Foundation of China under Grant 82261138629.
Appendix
Abstract. In this supplementary material, we provide additional information about the proposed method. Appendix A illustrates ablation studies on AdaMatch-Cyclic. Appendix B demonstrates visual results of CXR-to-report generation, report-to-CXR generation, and adaptive patches. Appendix C provides the implementation details of the proposed method.
Appendix A Ablation Studies on AdaMatch-Cyclic
@@
Table 9: The effectiveness of keywords and keypatches. B-4, M, and R-L represent BLEU-4, METEOR, and ROUGE-L, respectively.
Keywords Keypatches B-4 M R-L FID
@@
Table 10: The analysis on the different number of keywords ($N_{w}$). B, M, and R-L represent BLEU, METEOR, and ROUGE-L, respectively.
$N_{w}$ B-1 B-2 B-3 B-4 M R-L
@@
Figure 11: Visualization of texts and the corresponding adaptive patches. The boxes and arrows in different colors show adaptive patches and their center shifts from fixed patches.
Appendix B Visual Results
@@
CXR-to-Report Generation. As shown in Fig. 9, we compare CXR-to-report generation performance with existing methods on the MIMIC-CXR and OpenI datasets. In the first row, the proposed AdaMatch-Cyclic can capture the ‘Feeding tube tip’ and ‘mild bibasilar atelectasis’, while M2KT Yang et al. (2023) cannot observe such device and lung abnormality. The second case of the OpenI dataset shows that ‘cardiac silhouette is mildly enlarged’ and ‘degenerative changes of the thoracic spine’ can be discovered by our method. These imply that our AdaMatch-Cyclic can generate a more comprehensive and complete medical report with the guidance of generated keywords in comparison to current methods.
@@
Appendix C Implementation Details
In the AdaMatch-Cyclic model, we first train the AdaMatch model and then use the frozen AdaMatch to train LLM. The AdaMatch model consists of an image encoder and a text encoder. We utilize the DPT-medium Chen et al. (2021b) as the image encoder that includes AdaPatch module in stage 2 and 3. The image encoder is pre-trained on the MIMIC-CXR dataset with the disease classification task. We adopt the pre-trained BioClinicalBERT Alsentzer et al. (2019) as text encoder. The image and text encoders are followed by two convolutional layers with batch normalization and the ReLU activation function, respectively, to reduce the feature dimension to 256. The patch size $s$ is set to 4 for stage 1, and 2 for stage 2, 3, and 4. The number of sampled feature points $m$ is 3. We optimize the AdaMatch using LAMB optimizer ($\beta_{1}=0.9,\beta_{2}=0.999,\epsilon={10}^{-4}$). The cosine learning rate scheduler with the base learning rate ($lr\_b$) of $6\times 10^{-3}$ is adopted to linearly warm up to the peak learning rate (i.e. $lr\_p=lr\_b\times\sqrt{\frac{bs\_all}{512}}$) during the first quarter of the total training epochs, where $bs\_all$ denotes the effective total batch size. The total training epoch is 15 and the per GPU batch size is 112. In AdaMatch-Cyclic, we use the pre-trained VQ-GAN Esser et al. (2021) model to encode CXR images into image tokens and decode image tokens into CXR images. We adopt the dolly-v2-3b Conover et al. (2023) model as the pre-trained LLM. The LLM has 5,1845 token types with the first 5,0821 token types for text tokens and the rest 1,024 token types for image tokens. We add 5,0821 to each image token value encoded by VQ-GAN. To decode the image tokens into images, we subtract 5,0821 from the image tokens generated by LLM and feed image tokens to VQ-GAN decoder to obtain the generated CXR images. We train the LLM with the AdamW Loshchilov and Hutter (2018) optimizer. The learning rate is initialized as $5\times{10}^{-6}$ and the total training epoch is 5. The per GPU batch size is set to 24. The hyper-parameters $\kappa_{0},\kappa_{1},\kappa_{2},\kappa_{3},\kappa_{4}$ are set as 200, 10, 1000, 20, and 5, respectively. In CXR-to-report generation, we extract keywords for each adaptive patch and use the keywords with the top 10 patch-word similarities as hints of LLM. In addition, we use 5 keypatches to guide the LLM in report-to-CXR generation. All the experiments are conducted on 8 Nvidia A100 40GB GPUs. There is no overlap of patients among different subsets. We discard the medical reports with fewer than 3 tokens from both datasets. All the CXR images with different sizes are resized to $256\times 256$ pixels.